# Controlling noisy herds

### v2

@@ Abstract
Controlling multi-agent systems is a persistent challenge in organismal, robotic, and social collectives, especially when agents exhibit stochastic indecisiveness – frequently switching between conflicting behavioral rules. Here, we investigate the control of such noisy indecisive collectives through the lens of century-old sheepdog trials, where small groups of sheep exhibit unpredictable switching between fleeing and following behaviors. Unlike cohesive large flocks, these small indecisive groups are difficult to control, yet skilled dog-handler teams excel at both herding and precisely splitting them (shedding) on demand. Using a stochastic model, we introduce two central parameters – pressure (stimulus intensity) and lightness (response isotropy) – to simulate and quantify herding and shedding dynamics. Light sheep rapidly reach stable herding states, while heavy sheep exhibit intermittent herding and orthogonal alignment to the dog. High response isotropy fosters group cohesion but complicates splitting tasks, highlighting the nuanced trade-offs in collective control of noisy herds. Surprisingly, we find that stochastic indecisiveness, typically perceived as a challenge, can be leveraged as a critical tool for efficient control, enabling controlled herding and splitting of noisy groups. Building on these insights, we develop the Indecisive Swarm Algorithm (ISA) for artificial agents and benchmark its performance against standard algorithms, including the Averaging-Based Swarm Algorithm (ASA) and the Leader-Follower Swarm Algorithm (LFSA). ISA minimizes control energy in trajectory-following tasks, outperforming alternatives under noisy conditions. By framing these results within a stochastic temporal network framework, we show that even with a probabilistic description of the future dynamics, network restructuring (temporality) enhances control efficiency in a specific class of control problems. These insights establish a scalable framework for controlling noisy, behavior-switching collectives, with applications in swarm robotics, cellular engineering, opinion dynamics, and temporal networks.
Control challenges in behavior-switching collectives
@@
Emergent collective dynamics, where simple local interactions give rise to complex global behaviors, govern a wide range of systems. Examples include swarm robotics [1], animal collectives [2], social networks such as opinion dynamics [3], pedestrians’ movements [4], and vehicular traffic [5]. Controlling these systems poses significant challenges, as their behaviors often defy traditional control methods [6, 7, 8]. Unlike systems with predictable, linear dynamics, emergent systems require multiscale strategies that address both the microscopic interactions between individual agents and the macroscopic patterns that emerge at the group level [8].
@@
Biological collectives exemplify these challenges and introduce additional complexity when individuals switch between behaviors unpredictably in response to environmental factors [9, 10, 11, 12, 13] (see SI Section 1 and Table S1 for a full list of behavior-switching systems from ants and locusts to seals and even humans). Schools of golden shiners (a type of fish) stochastically transition between swarming, schooling, and milling [14]; carrier ants transporting cargo alternate roles between lifters and pullers based on their orientation and the nest’s position [9], and sheep in small flocks randomly switch between leading and following roles [12]. These systems highlight the need for robust control strategies that account for the stochastic and context-dependent nature of individual behavior transitions and their cascading effects on group dynamics. Despite advances in control theory and swarm robotics [7, 8], managing these dynamics remains an open challenge, particularly in systems where behavioral switching transitions are inherently noisy.
@@
Predator-prey systems provide a natural framework for studying the challenges of controlling noisy collectives with behavior-switching dynamics. For instance, flocks of starlings confuse raptors by transitioning between complex dynamical patterns. Similarly, large herds of wildebeests intermittently shift between selfish herding and solitary flight when confronted by predators like cheetahs. In response, predators adopt simplified strategies like focusing on a fixed point in space rather than tracking individual prey [15, 16] and split vulnerable individuals before leveraging speed and agility to secure their target [17, 18, 19]. These examples suggest that effective control of behavior-switching collectives does not always require precise tracking or prediction of behavioral transitions.
Unlike large, uncontrolled herds or purely theoretical systems, small flocks of sheep in competitions called sheep-dog trials provide a controlled environment where behavior-switching dynamics can be observed, quantified and systematically analyzed (see SI Section 2&3 for history and competition rules). At these trials, under threat, sheep oscillate stochastically between fleeing and regrouping, generating noisy, indecisive behaviors that challenge handlers and their dogs to cohesively herd or split groups under dynamic conditions (SI Video 1, 2). While prior studies have explored behavior-switching dynamics in the absence of a stimulus, in this work we uniquely frame sheep-dog trials as a control problem for "indecisive collectives" — systems where agents stochastically alternate between conflicting behaviors in presence of an external agent.
By introducing the metrics of "pressure" (stimulus intensity) and "lightness" (response isotropy), we bridge empirical observations with quantitative modeling to analyze various tasks in the sheepdog trials competition. Our findings reveal that indecision, often perceived as inefficiency, can paradoxically enhance control under specific conditions. Extending this framework, we propose a general Indecisive Swarm Algorithm (ISA) for artificial systems and demonstrate its effectiveness in trajectory-following tasks. By situating sheep-dog trials within the broader context of stochastic temporal networks, this work offers foundational principles for controlling noisy, behavior-switching collectives in applications ranging from swarm robotics to collective decision-making.
Sheep-dog Trials
@@
Historically, shepherds have exploited predator-prey dynamics to control collectives, leveraging herding dogs to manage farm animals as early as 1700 BC (see Figure  1a, SI Section 2) [20]. When a solitary sheep encounters a threat, it flees; in large groups, sheep exhibit selfish herd behavior [21]. However, in small groups, sheep struggle to choose a survival strategy, indecisively switching between solitary and collective behaviors, creating unpredictability (Figure  1b and SI Video 1). This unpredictability inspired the creation of sheep-dog trials, a 100-year-old sport testing a dog’s ability to control small sheep groups ($N_{s}\leq 5$[22]. In these trials, handlers and dogs not only move sheep cohesively (called herding) (Figure  1c-e), but also split the groups into subgroups (called shedding) (Figure  1f-h), showcasing the dog’s skill in managing indecisive collectives (see SI Section 3)  [23].
@@
Building on this rich history, we systematically investigate the dynamics of sheep-dog interactions to understand how external control influences indecisive collectives. Sheep-dog trials offer a wealth of qualitative knowledge, developed over generations by dog trainers and shepherds, about these interactions. Terms such as "pressure" and "lightness" capture key aspects: pressure refers to the threat perceived by sheep from a dog’s actions, such as approaching, barking, or staring, while lightness describes the sheep’s responsiveness to these cues [24]. Trained dogs apply pressure to herd or shed (split) sheep, and handlers also contribute by exerting pressure through their body posture during shedding. Lightness is a measure of the responsiveness of the sheep. Light sheep respond to minimal pressure but may panic under high pressure, whereas heavy sheep resist until high pressure is applied directly from the front. Assessing sheep lightness early in trials is essential for achieving effective control [24].
@@
To simplify analysis and exploit rotational symmetry, without loss of generality, we fix the reference frame such that the external stimulus (dog) is always positioned in the South (S) direction as a convention. To compare our simulation with the experimental data, we bin the orientations of the simulated agents into 4 possible directions: North (N), South (S), East (E), and West (W) (Figure 2a). Unlike conventional Vicsek-type models, where agents average the influences from neighbors and stimuli at each time step (from now on referred to as averaging agents[26, 27], our approach allows agents to be influenced by one factor at a time, stochastically switching between them. This design builds on stochastic choice models that are widely used to capture stochastic switching dynamics in diverse systems across scales ranging from cancer cells and insects to fish schools and human opinion dynamics [28, 11, 29, 30, 31]. For sheep, this method incorporates prior evidence that small flock decision-making is Markovian, with frequent switches between leader and follower roles, even in the absence of external stimuli [32, 12].
@@
To capture the time evolution of these transitions, we model the system as a stochastic process governed by a master equation. The master equation is widely used across fields like reaction kinetics, population dynamics, and network theory to describe how probabilities of different system micro-states evolves over time [33, 34, 35, 36]. For sheep orientation dynamics, it captures the interplay of stochastic influences, social interactions, and external stimuli driving collective behaivor.
@@
To compare the controllability of the two models, we simulate the herding process under different noise levels: $\epsilon/\gamma=0.08$ (Figure 4a, top) and $\epsilon/\gamma=0.8$ (Figure 4a, bottom). Starting from a random initial orientation, we find that at low relative noise, averaging agents outperform indecisive agents, reaching the target faster (Figure 4a, “initial" and “final"). However, at high relative noise, averaging agents fail to reach the target, while indecisive agents, although slower, successfully complete the task. In shedding tasks, averaging agents fail to split in two subgroups under all noise levels, whereas, indecisive agents consistently succeed, regardless of the magnitude of relative noise (Figure 4b).
@@
Exploiting Indecisiveness for Better Control
Having demonstrated that indecisiveness can aid control, we extend this concept to artificial systems for broader applications. Specifically, we investigate whether stochastic indecisiveness can improve control strategies in robotics, particularly for multi-agent systems navigating constrained trajectories. To test this, we consider the classic trajectory-following problem, a widely studied challenge in robotics [41, 42, 43]. We model a swarm of agents without visual sensing (blind agents) that rely solely on communicated orientations from other agents. A controller agent with visual sensing capability is tasked to steer the swarm from an initial position to a target position along a predefined trajectory by applying repulsion forces on individual agents.
@@
For LFSA, fixed pairwise non-reciprocal interactions often cause the swarm to split into small clusters, preventing it from reaching the target (SI video 11). ASA agents, on the other hand, exhibit strong averaging behavior (strong ordering), which makes them prone to overshooting and alignment challenges at the trajectory boundaries. This requires the controller to apply high input intensities to reorient the agents in the preferred direction, resulting in peaks in $I_{min}(t)$ ISA agents, however, reduce the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents. As a result, ISA agents successfully reach the target with significantly lower $I_{min}(t)$ compared to ASA.
@@
Figure 6: Indecisive Swarm Algorithm: a A schematic illustrates how temporality $\mathcal{T}$ influences a swarm of agents steered by a controller. Each snapshot depicts the agents’ orientation, the controller’s relative position, and their interactions. The network restructures every $\tau_{n}$, and agents update their directions at every dynamic update $\tau_{d}$ by considering interactions occurring between consecutive dynamic updates. When $\mathcal{T}=3$, the network restructures three times between two dynamics updates, and agents calculate a weighted average of all interactions at each dynamic update to determine their new direction. As $\mathcal{T}\rightarrow\infty$, this behavior resembles an averaging swarm algorithm (ASA), where agents average inputs from all the others and the controller’s repulsion to update their direction. In contrast, when $\mathcal{T}=0.5$, the network updates every two dynamic updates. As $\mathcal{T}\rightarrow 0$, the system mimics a leader-follower swarm algorithm (LFSA), where agents randomly follow a chosen agent or respond to the controller, indefinitely. At $\mathcal{T}=1$, the system operates as an indecisive swarm algorithm (ISA), with each network update directly followed by a dynamic update. b Trajectories of LFSA, ISA, and ASA agents are shown for different stimulus intensities $I$ with $N_{s}=50$. At high stimulus intensity, all agents follow the predefined path. At low stimulus intensity, only ISA agents stay on the predefined path, while ASA and LFSA agents deviate significantly. c-Top. The constrained predefined path used for control energy $\mathcal{E}$ calculations. ISA and ASA agents successfully reach the target, while LFSA agents fail. c-Bottom $I_{min}(t)$ for ISA and ASA agents. The peak in $I_{min}$ for ASA corresponds to the moment when the swarm reaches the boundary. d $\mathcal{E}$ and failure fraction shown as a functions of $\mathcal{T}$ for 1,500 simulations. As $\mathcal{T}<1$, the failure fraction increases for large group sizes ($N_{s}=50$). Control energy $\mathcal{E}$ achieves a minimum at $\mathcal{T}=1$, demonstrating the optimality of ISA for herding noisy agents along predefined trajectories.

### v3

# Temporal network restructuring improves control of indecisive collectives

@@ Abstract
Controlling stochastic temporal networks remains an open challenge in control theory. While predictable temporal networks with known future dynamics enhance controllability, real-world networks often exhibit stochasticity and unpredictability, making control harder. Here, we investigate control mechanisms for stochastic temporal networks by analyzing how biological controllers, such as shepherd dogs, manage panicked flocks of sheep. We studied a century-old shepherding competition, the sheepdog trials, where small groups of sheep unpredictably switch between fleeing and following behaviors–effectively forming stochastic temporal networks. Unlike large, cohesive flocks, these small, indecisive flocks are difficult to control, yet skilled dog-handler teams excel at both herding and splitting them (shedding) on demand. Using a stochastic choice model to describe the sheep’s behavioral shifts, we found that trained dogs exploit stochastic indecisiveness, typically seen as an obstacle, as a control tool, enabling both herding and splitting of noisy groups of sheep. Building on these insights, we developed the Indecisive Swarm Algorithm (ISA) for artificial agents and benchmarked its performance against standard approaches, including the Averaging-Based Swarm Algorithm (ASA) and the Leader-Follower Swarm Algorithm (LFSA). ISA minimizes control energy in trajectory-following tasks and outperforms alternatives under noisy conditions. By framing these results within a stochastic temporal network perspective, we demonstrate that even probabilistic knowledge of future dynamics can enhance control efficiency in specific scenarios. These findings establish a framework for managing stochastic temporal networks with applications in noisy, behavior-switching animal collectives, swarm robotics, and opinion dynamics.
Introduction
@@
Emergent collective dynamics, where simple local interactions give rise to complex global behaviors, govern a wide range of systems. Examples include swarm robotics [1], animal collectives [2], social networks such as opinion dynamics [3], pedestrians’ movements [4], and vehicular traffic [5]. Controlling these systems is challenging, as their behaviors often defy traditional control methods [6, 7, 8]. Unlike systems with predictable, linear dynamics, emergent systems are best described as complex networks that require multiscale strategies to address both the microscopic interactions between individual agents and the macroscopic patterns that emerge at the group level [8].
@@
Most of these networked systems introduce additional complexity when individual agents (nodes) switch between different behaviors, leading to temporal restructuring in the network. Biological collectives and social interactions in humans serve as prime examples of such behavior switching  [9, 10, 11, 12, 13] (see SI Section 1 and Table S1 for a full list of behavior-switching systems from ants and locusts to seals and humans). Carrier ants transporting cargo alternate roles between lifters and pullers based on their orientation and the nest’s position [9], sheep in small flocks randomly switch between leading and following roles [12], and during an epidemic outbreak, humans frequently switch between different interaction partners, facilitating spread of diseases [14]. These systems highlight the need for control strategies that account for the stochastic and context-dependent nature of individual behavior transitions and their cascading effects on the evolving temporal networks, where edges dynamically reorganize over time [15, 16, 17].
Recently, it has been shown that temporal restructuring can improve the controllability of a network [18]. Specifically, temporal networks require less time and less energetic cost to be controlled than their static counterparts [19, 20]. This counterintuitive observation relies on the fact that the future dynamics of the network are predictable and are exploited in designing the controls in the previous steps. However, when the switching dynamics are stochastic and unpredictable, temporality can make the control process more energetically demanding compared to a static network [21]. Therefore, despite advances in control theory and swarm robotics [7, 8], managing the dynamics of stochastic temporal networks remains an open challenge, particularly in systems where individual agents exhibit behavior switching.
@@
Predator-prey systems provide a natural framework for studying the challenges of controlling such noisy networks with behavior-switching dynamics. For instance, flocks of starlings confuse raptors by transitioning between complex dynamic patterns. Similarly, large herds of wildebeests intermittently shift between selfish herding and solitary flight when confronted by predators like cheetahs. In response, predators, instead of complex control mechanisms, adopt simplified strategies like focusing on a fixed point in space rather than tracking individual prey [22, 23]. This allows them to split vulnerable individuals before leveraging speed and agility to secure their target [24, 25, 26]. These examples suggest that effective control of stochastic temporal networks with behavior-switching individuals does not always require precise prediction of behavioral transitions.
In this work, we analyze such control mechanisms by studying shepherd dogs managing small flocks of sheep in a competition called the sheep-dog trials. Two key features of these competitions make them model systems for investigating control mechanisms in stochastic temporal networks. First, during these trials, when threatened, panicked sheep oscillate unpredictably and indecisively between fleeing from the dog and following other sheep, forming a stochastic temporal network. Trained shepherd dogs are highly effective at managing these noisy flocks under fluctuating conditions (SI Video 1). Second, unlike interactions between predators and large herds of animals in the wild, the sheepdog trials competitions provide a controlled environment where the behavior-switching dynamics of the sheep can be observed, quantified, and analyzed (see SI Section 2&3 for history and competition rules).
By bridging empirical observations with quantitative modeling to analyze various tasks in the sheepdog trials competition, we find that shepherd dogs utilize the behavior switching in sheep for herding and splitting (shedding) the flocks. Behavior switching dynamics have been previously studied in the context of animal collectives and human societies (voter models) using individual-based stochastic choice models [27, 11, 28, 29, 30]. In this work, we build on the existing framework to frame sheep-dog trials as a control problem for "indecisive collectives" — systems where agents stochastically alternate between different behaviors and interaction partners in the presence of an external agent.
This paper is structured as follows: We begin by exploring the nuances and rules of sheepdog trials. Next, we present a stochastic framework to develop quantitative metrics such as "pressure" and "lightness" that capture the nuanced behavior of sheep. The framework is based on qualitative insights from experienced handlers, and empirical data on sheep-dog dynamics. We then present a stochastic choice model and the master equation to model the indecisive transitions in sheep movement, comparing our model’s predictions with observed dynamics. Building on this, we investigate whether sheep indecisiveness could benefit the dog. Our findings reveal that stochastic indecisiveness can aid the dog in both herding and shedding tasks. Finally, we extend our analysis to develop the Indecisive Swarm Algorithm (ISA), a swarm control strategy inspired by shepherding dynamics. By modeling ISA as a non-reciprocal stochastic temporal network and comparing it against the standard Averaging-Based Algorithm (ASA) and Leader-Follower Swarm Algorithm (LFSA), we demonstrate that for specific control tasks like herding, ISA minimizes control energy requirements.
Results
Terminology in Sheep-dog Trials
@@ Results
Historically, shepherds have exploited predator-prey dynamics to control collectives, leveraging herding dogs to manage farm animals as early as 1700 BC (see Figure  1a, SI Section 2) [31]. When a solitary sheep encounters a threat, it flees; in large groups, sheep exhibit selfish herd behavior [32]. However, in small groups, sheep struggle to choose a survival strategy, indecisively switching between solitary and collective behaviors, creating unpredictability (Figure  1b and SI Video 1). This unpredictability led to the creation of sheep-dog trials, a 100-year-old sporting competition testing a dog’s ability to control small sheep groups ($N_{s}\leq 5$[33]. In these trials, handlers and dogs not only move sheep cohesively (called herding), but also split the groups into subgroups (called shedding), showcasing the dog’s skill in managing indecisive collectives (Figure  1c-h, see SI Section 3)  [34].
@@
In sheep-dog trials, qualitative terms such as "pressure" and "lightness" convey the following aspects: pressure refers to the threat perceived by sheep from a dog’s actions, such as approaching, barking, or staring, while lightness describes the sheep’s responsiveness to these cues [35]. Trained dogs apply pressure to herd or shed (split) sheep, and handlers also contribute by exerting pressure through their body posture during shedding. Lightness is a measure of the responsiveness of the sheep. Light sheep respond to minimal pressure but may panic under high pressure, whereas heavy sheep resist until high pressure is applied directly from the front. Assessing sheep lightness early in trials is essential for achieving effective control [35].
@@
To simplify analysis and exploit rotational symmetry, without loss of generality, we fix the reference frame such that the external stimulus (dog) is always positioned in the South (S) direction as a convention. To compare our simulation with the experimental data, we bin the orientations of the simulated agents into 4 possible directions: North (N), South (S), East (E), and West (W) (Figure 2a).
Traditionally, the dynamics of sheep are modeled with a conventional averaging-based approach, where each sheep averages the orientations of its neighbors and threat from the dog before deciding the direction of motion at each time step (from now on referred to as averaging agents)  [37, 38]. While this fully connected static network approach captures the dynamics of large flocks of sheep, it fails to simulate the indecisive behavior-switching dynamics in small flocks. Our approach allows agents to be influenced by one factor at a time, stochastically switching between them. This design is an extension of stochastic choice models that are widely used to capture stochastic switching dynamics in diverse systems across scales ranging from cancer cells and insects to fish schools and human opinion dynamics (such as two state voter models) [27, 11, 28, 29, 30]. For sheep, this method incorporates prior evidence that small flock decision-making is Markovian, with frequent switches between leader and follower roles, even in the absence of external stimuli [39, 12].
@@
To capture the time evolution of these transitions, we model the system as a stochastic process governed by a master equation. The master equation is widely used across fields like reaction kinetics, population dynamics, and network theory to describe how probabilities of different system micro-states evolve over time [40, 41, 42, 43]. For sheep orientation dynamics, it captures the interplay of stochastic influences, social interactions, and external stimuli driving collective behaivor.
@@
To compare the controllability of the two models, we simulate the herding process under different noise levels: $\epsilon/\gamma=0.08$ (Figure 4a, top) and $\epsilon/\gamma=0.8$ (Figure 4a, bottom). Starting from a random initial orientation, we find that at low relative noise, averaging agents outperform indecisive agents, reaching the target faster (Figure 4a, “initial" and “final"). However, at high relative noise, averaging agents fail to reach the target, while indecisive agents, although slower, successfully complete the task. In shedding tasks, averaging agents fail to split into two subgroups under all noise levels, whereas, indecisive agents consistently succeed, regardless of the magnitude of relative noise (Figure 4b).
@@
Developing an Indecisive Swarm Algorithm (ISA)
Having demonstrated that indecisiveness can improve control, we extend this concept to artificial systems for broader applications. Specifically, we investigate whether stochastic indecisiveness can improve control strategies in robotics, particularly for multi-agent networks navigating constrained trajectories. Control mechanisms that are effective for single agents often fail in multi-agent networked systems due to the emergence of complex global behaviors from simple local interactions between agents. Designing controllers capable of simultaneously managing local, microscopic interactions and global network-level dynamics is still an open problem in control theory [8, 48]. We show that by introducing stochastic indecisiveness in the multi-agent system, one can enable simple controllers, designed to control single agents, to control multi-agent networked systems.
To illustrate this, we use the classic trajectory-following problem, a widely studied challenge in robotics [49, 50, 51]. We model a swarm of agents without visual sensing (blind agents) that rely solely on communicated orientations from other agents. A controller agent with visual sensing capability is tasked to steer the swarm from an initial position to a target position along a predefined trajectory by applying repulsion forces on individual agents. The problem is a special class of controllability problems called herdability [52, 53, 54].
@@
The optimality of ISA can be attributed to its ability to combine the best features of ASA and LFSA. ASA agents quickly align with each other due to their averaging behavior, but they exhibit a strong order that requires the controller to apply high input intensities to reorient the agents in the preferred direction (Figure 7a) . For LFSA, due to the network hierarchy, the information from the controller can reach the follower nodes even if the input signal strength is low. However, fixed pairwise non-reciprocal interactions often cause the swarm to split into small clusters, preventing it from reaching the target (Figure 7b, SI video 11). ISA reduces the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents. As a result, ISA agents successfully reach the target with significantly lower $I_{min}(t)$ compared to ASA (Figure 7c).
Figure 6: Indecisive Swarm Algorithm. a A schematic illustrates how temporality $\mathcal{T}$ influences a swarm of agents steered by a controller. Each snapshot depicts the agents’ orientation, the controller’s relative position, and their interactions. The network restructures every $\tau_{n}$, and agents update their directions at every dynamic update $\tau_{d}$ by considering interactions occurring between consecutive dynamic updates. When $\mathcal{T}=3$, the network restructures three times between two dynamics updates, and agents calculate a weighted average of all interactions at each dynamic update to determine their new direction. As $\mathcal{T}\rightarrow\infty$, this behavior resembles an averaging swarm algorithm (ASA), where agents average inputs from all the others and the controller’s repulsion to update their direction. In contrast, when $\mathcal{T}=0.5$, the network updates every two dynamic updates. As $\mathcal{T}\rightarrow 0$, the system mimics a leader-follower swarm algorithm (LFSA), where agents randomly follow a chosen agent or respond to the controller, indefinitely. At $\mathcal{T}=1$, the system operates as an indecisive swarm algorithm (ISA), with each network update directly followed by a dynamic update. b Trajectories of LFSA, ISA, and ASA agents are shown for different stimulus intensities $I$ with $N_{s}=50$. At high stimulus intensity, all agents follow the predefined path. At low stimulus intensity, only ISA agents stay on the predefined path, while ASA and LFSA agents deviate significantly. c-Top. The constrained predefined path used for control energy $\mathcal{E}$ calculations. ISA and ASA agents successfully reach the target, while LFSA agents fail. c-Bottom $I_{min}(t)$ for ISA and ASA agents. The peak in $I_{min}$ for ASA corresponds to the moment when the swarm reaches the boundary. d $\mathcal{E}$ and failure fraction shown as a functions of $\mathcal{T}$ for 1,500 simulations. As $\mathcal{T}<1$, the failure fraction increases for large group sizes ($N_{s}=50$). Control energy $\mathcal{E}$ achieves a minimum at $\mathcal{T}=1$, demonstrating the optimality of ISA for herding noisy agents along predefined trajectories.
Figure 7: Schematic network representation of ASA, LFSA and ISA. The red node (node 1) represents the controller. a In ASA, due to the network being fully connected, a high input strength is required from the controller to align the agents in the preferred orientation. b In LFSA, due to the fixed pairwise non-reciprocal interactions, the network often splits in small clusters making it impossible to control. c ISA reduces the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents.